OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 ANSWERS : –x – 5y – 6 x + 5y – 6 –x – 5y + 6 x + 5y + 6

OpenStudy (amistre64):

its easier to me to just add z and subtract 6 from the first eq ...

OpenStudy (anonymous):

ADD z to what ?

OpenStudy (amistre64):

x + 5y – z = 6 +z +z ------------- x + 5y + 0 = 6 + z -6 -6 ---------------- x + 5y -6 = z

OpenStudy (amistre64):

add it to both sides of the equation ...